Chapter 12 H.W. Begin Date: I 00.00 AMDe Date 00 PM End Date (20%) Problem 5: Hoover Dam on the Colorado River nde baghest dn.n the United States at 221m, with a power output of 680 MW The dam generates electricity by flowing water down to a poust 150 m below the stop, at an average flow rate of 65o ▲ 50% Part (a) Calculate the power in das flow watts. Grade Samary Power Pnmal 100% sino cos) asino acos tano Degreesads Fedback: dedactaioe pe ndback ▲ 50% Part (b) w aat is the titio of das power to the facility s aetage of 680 MW?
Answer #1

Assume all PE of the water is converted to electrical power.

P = Work/s = mgh/s = (650m^3)(1000kg/m^3)(9.8N/kg)(150m)/1s =9.56 *10^8 W

Ratio = 9.56 *10^8 W/680 MW= 1.4